Find the length of the indicated side of triangle PQR.
1. The measure of Q= 90, measure of P=53 QR= 197, PR=? (find what PR equals)
I just need an equation to solve this using trig.

Answer by scott8148(6628)

since Q=90, this is a right triangle with PR as the hypotenuse and QR the side opposite angle P

using trig, sin P=QR/PR ... so PR=QR/(sin P) ... PR=197/(sin 53)